Directions: Solve the equation by completing the square. Express any complex numbers using i notation.
2x^2+4x+1=0

Answer by fractalier(6550)

Okay from
2x^2 + 4x + 1 = 0
the first thing we do is divide by 2 and get
x^2 + 2x + 1/2 = 0
then complete the square
x^2 + 2x + 1 = 1/2
(x + 1)^2 = 1/2
now take the root
x + 1 = ± sqrt(1/2)
x = -1 ± sqrt(1/2)
x = -1 ± [sqrt(2) / 2]